THE SLOPE FORM OF THE EQUATION OF THE LINE PERPENDICULAR TO THE GRAPH OF Y = -3/2x - 7 THAT PASSES THROUGH (3,-2).
Answer by Alan3354(69443)

PERPENDICULAR TO THE GRAPH OF Y = -3/2x - 7 THAT PASSES THROUGH (3,-2).
---------
The slope of y = (-3/2)x - 7 is -3/2.
The slope of lines perpendicular is the negative inverse, = 2/3
Use y = mx + b and the point to find b
-2 = (2/3)*3 + b
b = -4
---------
y = (2/3)x - 4